OpenStudy (anonymous):

2/3(c-18)=7 help me find c

OpenStudy (anonymous):

Multiply both sides by 3: 3 * 2/3(c-18) = 3* 7 2(c - 18) = 21 then divide both sides by 2: 2(c-18) / 2 = 21/2 c - 18 = 21/2 THen add 18 to both sides: c - 18 + 18 = 21/2 + 18 c = 21/2 + 18 Then just simplify the value of c by replacing 18 by 36/2: c = 21/2 + 18 = 21/2 + 36/2 = (21+36)/2 = 57/2 So c = 57/2 (or 23 1/2)

OpenStudy (anonymous):

thank you very much :)